# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 8.2. The ring of cooperations
One may now use the ideas of Remark 4.8 to study the kq-resolution one filtration at a time. The $n$-line of the kq-resolution may be computed by a motivic Adams spectral sequence taking the form
$E_{2}=\bigoplus_{I\in\mathcal{I}_{n}}\Sigma^{4|I|,2|I|}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(I))\implies\pi_{*+n,*}^{\mathbb{R}}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}).$
This spectral sequences collapses for the same reasons given above. Note that although there is an algebra structure on the $E_{1}$-page, the individual $n$-lines do not carry any ring structure, as $\text{kq}\otimes\overline{\text{kq}}^{\otimes n}$ is not a ring spectrum. These computation do present the $n$-line as a module over the 0-line, given by $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}}).$
Determining the differentials in the kq-resolution is much more difficult. By complex Betti realization, we can deduce much of the behavior of the complex kq-resolution from the bo-resolution (see Corollary 2.4). However, there is information in the real kq-resolution which is both $\rho$-torsion and $v_{1}$-periodic that is not detected by these methods. We plan to analyze the real kq-resolution in future work using a combination of these methods and $C_{2}$-equivariant homotopy theory.
@@ Appendix A Charts
In this appendix, we record charts depicting the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules described in Section 6. A black $\blacksquare$ represents $\mathbb{F}_{2}[\tau^{4}][\rho,h_{1}]$. A black $\bullet$ represents $\mathbb{F}_{2}[\tau^{4}]$. A vertical black line represents multiplication by $h_{0}$. A horizontal black line represents multiplication by $\rho$. A diagonal black line represents multiplication by $h_{1}$.
Figure 28. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $P$.
Figure 29. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $H$.
Figure 30. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $PH$.
Figure 31. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $D$.
Figure 32. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $S$.
Figure 33. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $T$.
Figure 34. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $J$.
Figure 35. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $JD$.

### v2

@@ 8.3. Application to the kq-resolution
Determining the differentials in the real kq-resolution seems to be quite difficult. By complex Betti realization, we can deduce much of the behavior of the complex kq-resolution from the bo-resolution (see Corollary 2.3). However, there is information in the real kq-resolution which is both $\rho$-torsion and $v_{1}$-periodic that is not detected by these methods. We plan to analyze the real kq-resolution in future work using a combination of these methods and $C_{2}$-equivariant homotopy theory.
@@ Appendix A Charts
In this appendix, we record charts depicting the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules described in Section 6. A black $\blacksquare$ represents $\mathbb{F}_{2}[\rho,\tau^{4}]$. A black $\bullet$ represents $\mathbb{F}_{2}[\tau^{4}]$. A vertical black line represents multiplication by $h_{0}$. A horizontal black line represents multiplication by $\rho$. A diagonal black line represents multiplication by $h_{1}$.
Figure A.1. The module $P$.
Figure A.2. The module $H$.
Figure A.3. The module PHPH.
Figure A.4. The module $D$.
Figure A.5. The module $S$.
Figure A.6. The module $T$.
Figure A.7. The module $J$.
Figure A.8. The module JDJD.